# Generalized Distribution Prediction for Asset Returns

### v1

@@ 4.3 Limitations and Challenges
While our approach yielded promising results, several limitations persist. One significant limitation is the inability to fairly compare our proposed approach with that of the authors in [2]. This issue arises primarily due to the lack of access to the original dataset used in their study and the absence of publicly available code for their implementation. Consequently, we were unable to replicate their exact methodology, which inherently limits the comparison between qLSTM approach and their dense model.
Furthermore, the two approaches differ fundamentally in how they process data. The qLSTM model is specifically designed to handle sequence data, allowing it to capture patterns across variable-length sequences and predict multiple quantiles for future time steps simultaneously. In contrast, the dense model used in [2] operates on fixed-length windows and focuses on single-step quantile predictions. This introduces a degree of inconsistency in the comparison, as the qLSTM’s ability to process sequential information may provide it with a natural advantage in capturing long-term dependencies in asset returns.
Another challenge emerged when comparing model performance on synthetic data. Due to differences in code constraints and hyperparameter configurations, the two models were not exposed to the same number of observations. Specifically, the qLSTM model used a normalization window of 34 days, while the qDense model operated on a longer window of 123 days. This discrepancy means that the qLSTM model was trained on a larger number of observations, increasing its potential for failure. However, despite being exposed to more opportunities for failure, the qLSTM model still outperformed the qDense model.
@@ 5 Conclusion
In this paper, we presented a novel model for predicting the distribution of asset returns using a quantile-based approach, extending prior research. By leveraging Long Short-Term Memory (LSTM) networks, we modeled the distribution of returns more dynamically and generalized these predictions across multiple asset classes, including synthetic data. Our approach diverged from previous models by incorporating asset-neutral features and utilizing market data to further refine the predictions.
The results demonstrated significant improvements over a baseline linear quantile regression model, outperforming it by 98% across all datasets. Additionally, the LSTM approach outperformed a dense network model by over 50% on both real-world and synthetic data. This performance boost underscores the strength of our approach in capturing the complexities of financial return distributions, particularly for high-volatility assets like cryptocurrencies. Furthermore, the model’s ability to maintain strong performance across diverse synthetic datasets highlights its generalizability and flexibility.
For future work, expanding the dataset to include more assets and asset classes—such as bonds, options, or credit default swaps—could further test the model’s generalizability. Additionally, exploring different time intervals, such as hourly, minute, or even second-level data, could provide deeper insights, as this study exclusively utilized daily intervals. Testing the model on illiquid assets could further validate its robustness in different market conditions. In addition, incorporating alternative data sources, such as social media sentiment, news feeds, or satellite imagery, might further enhance model accuracy. Finally, future research could focus on optimizing the LSTM-based model for real-world applications, such as integrating it into real-time trading systems or risk management frameworks.
Statements and Declarations
The authors declare that there are no competing interests related to this work.
Appendix A Financial Assets
@@
Bitcoin BTC-USD Cryptocurrencies 2014-10-09 2.25 59.09
Ethereum ETH-USD Cryptocurrencies 2017-12-01 1.16 76.15
Ripple XRP-USD Cryptocurrencies 2017-12-01 0.64 96.11
Cardano ADA-USD Cryptocurrencies 2017-12-01 1.14 94.15
Dogecoin DOGE-USD Cryptocurrencies 2017-12-01 2.69 115.10
Binance Coin BNB-USD Cryptocurrencies 2017-12-01 3.61 86.33
NEAR Protocol NEAR-USD Cryptocurrencies 2020-11-05 2.65 109.42
Bitcoin Cash BCH-USD Cryptocurrencies 2017-12-01 -1.20 93.81
Litecoin LTC-USD Cryptocurrencies 2014-10-09 1.38 85.60
Solana SOL-USD Cryptocurrencies 2020-05-02 5.85 114.08
Appendix B Market Data
@@
Appendix C qDense Training Details & Results
Hyperparameter Asset Model (Stage 1) Market Model (Stage 2)
Batch size $b\in\{32,64,128,256,512,1024,2048\}$
Learning rate $\eta\in[1\times 10^{-6},1\times 10^{-3}]$
Normalization window $w\in[5,250]$
Dropout rate $D\in[0.0,0.9]$
L1 regularization $\lambda_{1}\in[0.0,1\times 10^{-3}]$
L2 regularization $\lambda_{2}\in[0.0,1\times 10^{-3}]$
Use layer normalization Boolean: $\{1,0\}$
[…]
Hyperparameter Asset Model (Stage 1) Market Model (Stage 2)
Batch size 2048
Learning rate 0.0007
Normalization window 123
Raw hidden layers 2 4
Hidden layer units 64, 512 64, 512, 16, 32
Dropout rate 0.1780
Market activation Sigmoid
Hidden activation ELU
Use layer normalization Yes (1)
L1 regularization 0.0007
L2 regularization 0.0004
Appendix D Kernel Density Estimation
Kernel Density Estimation (KDE) is a non-parametric method used to estimate the probability density function (PDF) of a random variable. In this study, KDE was used to convert predicted quantiles into a smooth probability density function. The advantage of KDE is that it allows for flexible modeling of the underlying data distribution without assuming any particular parametric form, which is useful when working with financial return distributions that may not follow standard distributions like normal or log-normal.
The quantiles predicted by the model are transformed into a smooth PDF using KDE, ensuring that the final distribution is continuous and appropriately represents the data. This process helps provide more accurate density estimates, which are crucial for downstream tasks such as risk estimation and derivative pricing.
Below is the Python code used for the KDE conversion in this work.
1import numpy as np
2import matplotlib.pyplot as plt
3from sklearn.neighbors import KernelDensity
4
5def quantiles_to_pdf_kde(quantiles, grid_points=200, bandwidth=0.01):
6 """
7 Convert quantiles to a probability density function (PDF) using Kernel Density Estimation (KDE).
8
9 Parameters:
10 - quantiles: Array of quantile values.
11 - grid_points: Number of grid points for a denser representation.
12 - bandwidth: Bandwidth for the KDE (controls the smoothness of the density).
13
14 Returns:
15 - pdf_grid: Probability density function values on the dense grid.
16 - cdf_grid: Cumulative distribution function values on the dense grid (obtained by integrating the PDF).
17 - dense_grid: The dense grid of quantiles corresponding to the PDF and CDF values.
18 """
19
20 # Create a denser grid between the quantiles
21 dense_grid = np.linspace(min(quantiles), max(quantiles), grid_points).reshape(-1, 1)
22
23 # Fit KDE to the quantile values
24 kde = KernelDensity(kernel=’gaussian’, bandwidth=bandwidth).fit(quantiles.reshape(-1, 1))
25
26 # Compute log density and convert to regular density
27 log_dens = kde.score_samples(dense_grid)
28 pdf_grid = np.exp(log_dens)
29
30 # Approximate the CDF by cumulatively summing the PDF
31 cdf_grid = np.cumsum(pdf_grid)
32 cdf_grid /= cdf_grid[-1] # Normalize to [0, 1]
33
34 return pdf_grid, cdf_grid, dense_grid
Listing 1: Quantiles to PDF using KDE
Appendix E Synthetic Data
@@
Distribution Type Asset Generated $\mu$, $\sigma$ Correlation with Market Data
@@
1import numpy as np
2import pandas as pd
3from scipy.stats import gamma, lognorm, uniform
4
5def generate_synthetic_data_with_target_correlation(market_data, target_correlation, mu, sigma, distribution=’normal’):
6 """
7 Generate a synthetic dataset with returns that have a target correlation to the market_data.
8
9 Parameters:
10 - market_data: numpy array of log returns from the market data
11 - target_correlation: the target correlation for synthetic data’s log returns to the market data
12 - mu: mean of the synthetic dataset
13 - sigma: standard deviation of the synthetic dataset
14 - distribution: the distribution to sample synthetic data from (’normal’, ’gamma’, ’lognormal’, ’uniform’)
15
16 Returns:
17 - df: DataFrame with synthetic prices
18 """
19
20 num_periods = len(market_data)
21
22 # Step 1: Generate independent noise based on the specified distribution
23 if distribution == ’normal’:
24 noise = np.random.normal(loc=0, scale=1, size=num_periods) # Standard normal noise
25 elif distribution == ’gamma’:
26 shape = 2 # Shape parameter for the gamma distribution
27 noise = gamma.rvs(a=shape, scale=1, size=num_periods)
28 noise = (noise - np.mean(noise)) / np.std(noise) # Standardize noise
29 elif distribution == ’lognormal’:
30 noise = lognorm.rvs(s=1, scale=np.exp(0), size=num_periods)
31 noise = (noise - np.mean(noise)) / np.std(noise) # Standardize noise
32 elif distribution == ’uniform’:
33 noise = uniform.rvs(loc=-1, scale=2, size=num_periods) # Uniform distribution [-1, 1]
34 else:
35 raise ValueError("Unsupported distribution. Choose from ’normal’, ’gamma’, ’lognormal’, or ’uniform’.")
36
37 # Step 2: Standardize market data
38 market_std = (market_data - np.mean(market_data)) / np.std(market_data)
39
40 # Step 3: Create a linear combination of market data and noise for correlated returns
41 synthetic_log_returns = target_correlation * market_std + np.sqrt(1 - target_correlation**2) * noise
42
43 # Step 4: Adjust the synthetic returns to have the desired mean and standard deviation
44 synthetic_log_returns = synthetic_log_returns * sigma + mu
45
46 # Step 5: Convert synthetic log returns to price data
47 initial_price = np.random.uniform(5, 1000) # Random initial price
48 synthetic_prices = initial_price * np.exp(np.cumsum(synthetic_log_returns))
49
50 # Step 6: Create a DataFrame with synthetic price data
51 df = pd.DataFrame(synthetic_prices, columns=[’Adj Close’])
52
53 # Add High, Low, Volume columns to allow all functions to work
54 df[’High’] = df[’Adj Close’] * (1 + np.random.normal(0, 0.01, size=len(df)))
55 df[’Low’] = df[’Adj Close’] * (1 - np.random.normal(0, 0.01, size=len(df)))
56 df[’Volume’] = 0
57
58 df = generate_technical_features(df)
59
60 return df
@@ 16 Returns:
Appendix F Quantiles
The quantiles used in this study are:

### v2

@@ 4.3 Limitations and Challenges
While our approach yielded promising results, several limitations persist. One significant limitation is the inability to fairly compare our proposed approach with that of the authors in [12]. This issue arises primarily due to the lack of access to the original dataset used in their study and the absence of publicly available code for their implementation. Consequently, we were unable to replicate their exact methodology, which inherently limits the comparison between the qLSTM approach and their dense model.
In terms of model performance, while the qLSTM model demonstrated superior results in tail risk estimation and overall robustness, its slightly higher Value at Risk (VaR) at broader confidence levels (e.g., 0.05) compared to the Gaussian and qHybrid models suggests room for improvement in capturing broader distributional patterns. Additionally, the qHybrid model, while effective in several scenarios, exhibited slightly higher RMSE for moments like kurtosis, indicating potential difficulties in accurately modeling extreme tails for certain asset classes.
Finally, our results reveal some variability in model performance across asset classes. For example, while qLSTM excelled in cryptocurrencies and equity indices, the qDense and qHybrid models outperformed it in currency pairs and synthetic datasets with skewed distributions. These findings suggest that no single model is universally optimal, underscoring the importance of tailoring model architectures to specific data characteristics and use cases.
@@ 5 Conclusion
This paper introduces a novel approach to estimating the distribution of log returns by combining quantile regression with smoothed density estimation, providing a robust framework for capturing tail risks and distributional characteristics across diverse asset classes. By leveraging asset-neutral features and general market variables, our models are applicable to a wide range of financial datasets, making them versatile tools for risk management, portfolio optimization, and derivative pricing. The proposed qLSTM model consistently demonstrated superior performance on real-world data, achieving the lowest quantile loss across most asset classes, while the qHybrid model excelled in synthetic datasets, particularly for skewed or heavy-tailed distributions. These results underscore the strengths of both models in addressing the limitations of traditional distributional assumptions, such as normality, in financial returns.
The evaluation metrics, including Wasserstein distance, CRPS, and VaR, highlight the complementary nature of the qLSTM and qHybrid models. The qLSTM model excelled in extreme tail risk estimation, capturing critical distributional behaviors for volatile asset classes like cryptocurrencies and equity indices. Meanwhile, the qHybrid model provided improved calibration for broader confidence levels and distributions, demonstrating its adaptability to various market conditions. However, some variability in performance across asset classes suggests that no single model is universally optimal, emphasizing the need for tailored approaches based on specific data characteristics.
To enhance the literature, the paper allows for reproducibility by publicly hosting the code used for this study on GitHub. Furthermore, all data is sourced via an open API and the script to download the required data can be found in the GitHub repository. This makes comparisons and iterations on the works proposed, more accessible thereby fostering research robustness.
Future research can explore other methods to create better Hybrid models, particularly by leveraging certain quantiles with dynamic weighting schemes to optimize the combination of Gaussian and model-predicted quantiles. This could provide a more adaptable and context-specific qHybrid approach to estimating distributions. Moreover, extending the model to handle ultra-high-frequency data or incorporating more sophisticated market features, such as macroeconomic indicators or sentiment analysis, could further enhance its applicability. These avenues represent exciting opportunities to refine and expand the utility of distributional modeling in finance.
Appendix A Quantiles
The quantiles used in this study are:
Appendix B Quantile-Smooth Density Estimation
This method generates a smooth probability density function (PDF) from a set of quantiles and their associated probabilities, ensuring that the resulting representation adheres to the properties of a valid probability distribution. Starting with the quantiles, it removes duplicate or nearly identical points to maintain numerical stability, then creates a denser grid for better resolution. Using a monotonic cubic spline or fallback interpolation, it estimates the CDF, which is clamped to the range [0, 1] and adjusted to be strictly non-decreasing. The method rescales the CDF to start at 0 and end at 1, guaranteeing its validity as a cumulative function.
To derive the PDF, the method computes the CDF gradient, applying smoothing techniques such as convolution to remove spikes and ensure a well-behaved density curve. Non-negativity is enforced, and the PDF is normalized to integrate to one. The final output includes the smoothed PDF, the regenerated CDF, and the grid over which they are defined, offering a high-quality approximation of the distribution. Optional visualization helps validate the results, showing both the smoothed PDF and the corresponding CDF with respect to the input quantiles. This makes it a robust method for converting quantiles into a smooth and interpretable density function suitable for further analysis or modeling.
The function is based on three priors, $\epsilon$ which controls how close different quantiles are allowed to each other, min density which controls the minimum amount of density allowed in the PDF, and window which controls how many extreme the convolutions are for smoothing. The optimal priors are found via TPE on the validation data, similar to how the hyperparameters are found for the models themselves. Table 10 shows the best values found during optimization.
Model $\epsilon$ min density window
qLSTM 0.000253 0.027341 113
hybrid 0.000899 0.033704 43
Table 10: Optimal priors for the smooth PDF approximation
Below is the Python code used for the Quantile-Smooth Density Estimation conversion in this work.
Listing 1: Quantiles to PDF using Quantile-Smooth Density Estimation
def generate_smooth_pdf(quantiles, taus, min_density=1e-3, eps=1e-6, window=61):
␣␣␣␣GenerateasmoothedPDFfromquantileswithadditionalcontrolstopreventspikes
␣␣␣␣andensuretheCDFisbetween[0,1].
# Constants
GRID_POINTS = 1000
og_quants = quantiles.copy()
og_taus = taus.copy()
unique_mask = np.concatenate(([True], np.diff(quantiles) > eps))
quantiles = quantiles[unique_mask]
taus = taus[unique_mask]
# Create denser grid
grid_x = np.linspace(quantiles[0], quantiles[-1], GRID_POINTS)
# Monotonic spline for the CDF
try:
cdf_monotonic = PchipInterpolator(quantiles, taus, extrapolate=False)
cdf = cdf_monotonic(grid_x)
except Exception as e:
cdf = np.interp(grid_x, quantiles, taus)
# Clamp CDF to [0,1], then ensure it’s monotonically non-decreasing
cdf = np.clip(cdf, 0, 1)
cdf = np.maximum.accumulate(cdf)
# Rescale so that it starts exactly at 0 and ends exactly at 1
cdf -= cdf[0]
if cdf[-1] > 0:
cdf /= cdf[-1]
# Approximate PDF from finite differences (or use derivative if PCHIP)
density = np.gradient(cdf, grid_x)
smoothed_density = np.convolve(density, np.ones(window)/window, mode=’same’)
# Ensure non-negative and non-zero density
smoothed_density = np.maximum(smoothed_density, min_density)
# Normalize PDF to integrate to 1
area = np.trapz(smoothed_density, grid_x)
smoothed_density = smoothed_density / area
# regenerate CDF
cdf = np.cumsum(smoothed_density) * (grid_x[1] - grid_x[0])
return grid_x, smoothed_density, cdf
Appendix C Data
The following sections detail the data used for this study.
C.1 Financial Assets
@@
Bitcoin BTC-USD Crypto 2014-10-09 2.25 59.09
Ethereum ETH-USD Crypto 2017-12-01 1.16 76.15
Ripple XRP-USD Crypto 2017-12-01 0.64 96.11
Cardano ADA-USD Crypto 2017-12-01 1.14 94.15
Dogecoin DOGE-USD Crypto-currencies 2017-12-01 2.69 115.10
Binance Coin BNB-USD Crypto-currencies 2017-12-01 3.61 86.33
NEAR Protocol NEAR-USD Crypto-currencies 2020-11-05 2.65 109.42
Bitcoin Cash BCH-USD Crypto-currencies 2017-12-01 -1.20 93.81
Litecoin LTC-USD Crypto-currencies 2014-10-09 1.38 85.60
Solana SOL-USD Crypto-currencies 2020-05-02 5.85 114.08
C.2 Market Data
@@
C.3 Synthetic Data
@@
Dist Type Asset Generated $\mu$, $\sigma$ Correl -ation with Market Data
@@
import numpy as np
import pandas as pd
from scipy.stats
import gamma, lognorm, uniform
def generate_synthetic_data_with_target_correlation(market_data, target_correlation, mu, sigma, distribution=’normal’):
␣␣␣␣Generateasyntheticdatasetwithreturnsthathaveatargetcorrelationtothemarket_data.
␣␣␣␣Parameters:
␣␣␣␣-market_data:numpyarrayoflogreturnsfromthemarketdata
␣␣␣␣-target_correlation:thetargetcorrelationforsyntheticdata’slogreturnstothemarketdata
␣␣␣␣-mu:meanofthesyntheticdataset
␣␣␣␣-sigma:standarddeviationofthesyntheticdataset
␣␣␣␣-distribution:thedistributiontosamplesyntheticdatafrom(’normal’,’gamma’,’lognormal’,’uniform’)
␣␣␣␣Returns:
␣␣␣␣-df:DataFramewithsyntheticprices
num_periods = len(market_data)
# Step 1: Generate independent noise based on the specified distribution
if distribution == ’normal’:
noise = np.random.normal(loc=0, scale=1, size=num_periods) # Standard normal noise
elif distribution == ’gamma’:
shape = 2 # Shape parameter for the gamma distribution
noise = gamma.rvs(a=shape, scale=1, size=num_periods)
noise = (noise - np.mean(noise)) / np.std(noise) # Standardize noise
elif distribution == ’lognormal’:
noise = lognorm.rvs(s=1, scale=np.exp(0), size=num_periods)
noise = (noise - np.mean(noise)) / np.std(noise) # Standardize noise
elif distribution == ’uniform’:
noise = uniform.rvs(loc=-1, scale=2, size=num_periods) # Uniform distribution [-1, 1]
else:
raise ValueError(”Unsupporteddistribution.Choosefrom’normal’,’gamma’,’lognormal’,or’uniform’.”)
# Step 2: Standardize market data
market_std = (market_data - np.mean(market_data)) / np.std(market_data)
# Step 3: Create a linear combination of market data and noise for correlated returns
synthetic_log_returns = target_correlation * market_std + np.sqrt(1 - target_correlation**2) * noise
# Step 4: Adjust the synthetic returns to have the desired mean and standard deviation
synthetic_log_returns = synthetic_log_returns * sigma + mu
# Step 5: Convert synthetic log returns to price data
initial_price = np.random.uniform(5, 1000) # Random initial price
synthetic_prices = initial_price * np.exp(np.cumsum(synthetic_log_returns))
# Step 6: Create a DataFrame with synthetic price data
df = pd.DataFrame(synthetic_prices, columns=[’AdjClose’])
# Add High, Low, Volume columns to allow all functions to work
df[’High’] = df[’AdjClose’] * (1 + np.random.normal(0, 0.01, size=len(df)))
df[’Low’] = df[’AdjClose’] * (1 - np.random.normal(0, 0.01, size=len(df)))
df[’Volume’] = 0
df = generate_technical_features(df)
return df
@@
Appendix D Training Details
The following sections give details on training procedures including both training of the models as well as optimal hyper parameters.
D.1 qLSTM Hyperparamters
The following table shows the optimal hyperparameters for the qLSTM model.
Hyperparameter Asset Model (Stage 1) Market Model (Stage 2)
Batch size 64
Learning rate 0.0006
Normalization window 219
LSTM layers 1 1
LSTM units 16 16
Dense layers 5 3
Dense units per layer 128, 64, 64, 32, 32 16, 16, 32
Dropout rate 0.1778
Market activation Tanh
Hidden activation ELU
Use layer normalization no (0)
L1 regularization 0.0006
L2 regularization 0.0009
D.2 qDense Training Details & Results
Hyper-parameter Asset Model (Stage 1) Market Model (Stage 2)
Batch size $b\in\{32,64,128,256,512,1024,2048\}$
Learning rate $\eta\in[1\times 10^{-6},1\times 10^{-3}]$
Normalization window $w\in[5,250]$
Dropout rate $D\in[0.0,0.9]$
L1 regularization $\lambda_{1}\in[0.0,1\times 10^{-3}]$
L2 regularization $\lambda_{2}\in[0.0,1\times 10^{-3}]$
Use layer normalization Boolean: $\{1,0\}$
[…]
Hyperparameter Asset Model (Stage 1) Market Model (Stage 2)
Batch size 32
Learning rate 0.0003
Normalization window 207
Raw hidden layers 3 1
Hidden layer units 32, 32, 64 64
Dropout rate 0.4623
Market activation ELU
Hidden activation Leaky RELU
Use layer normalization No (0)
L1 regularization 0.0005
L2 regularization 0.0001